solve the system of equation by graphing.
9x-6y=42
6y-9x=-42

is it consistent or inconsistent
dependant or independant
i'm totally lost with this one, thinking there is no solution, but is that correct?
thanks

Answer by psbhowmick(878)

Let us consider the first equation: 9x - 6y = 42.
Multiplying both sides of this equation by (-1) we have
(-1)*(9x - 6y) = (-1)*42
(-1)*9x - (-1)*6y = -42
-9x + 6y = -42
i.e. 6y - 9x = -42
This is nothing but your second equation so the two equations are not inconsistent, but are the same equation instead and hence dependent.
If you try to draw the graphs, you will get ONLY one straight line as shown below.